# Unsupervised Evolutionary Cell Type Matching via Entropy-Minimized Optimal Transport

### v1

@@ 5.2 Cross-Species Correspondence of Bipolar Cell (BC) Types
We next applied OT-MESH to cross-species alignment of BC types between mouse (15 types) and macaque (12 types). The resulting correspondence matrix (Figure 3A) reveals clear evolutionary matches, forming strong diagonal-like patterns.
@@
Alignment scores (SupplementaryTable 1) quantify these correspondences. The strongest match was between mouse RBC and macaque RB (rod BCs), with a score of 0.91. Among cone BCs, macaque DB1, DB2, DB3a, DB3b, DB6, IMB, and BB/GB* types were uniquely and strongly assigned to mouse BC2, BC4, BC3A, BC3B, BC6, BC7, and BC8/9 types, respectively, consistent with previous reports hahnEvolutionNeuronalCell2023a ; shekharCOMPREHENSIVECLASSIFICATIONRETINAL2016a .
@@
5.3 Cross-Species Correspondence of Ganglion Cell (RGC) Types
@@
Most notably, our analysis confirmed the correspondence between mouse alpha RGCs and primate midget/parasol ganglion cells with high confidence (SupplementaryTable 2). Specifically, mouse OFF transient alpha (C45) maps strongly to macaque OFF parasol ganglion cells (PG-OFF) (score: 0.61), while ON sustained alpha (C43) corresponds to macaque ON midget ganglion cells (MG-ON) (score: 0.52). Similarly, mouse ON transient alpha (C41) and OFF sustained alpha (C42) map to macaque ON parasol (PG-ON) and OFF midget ganglion cells (MG-OFF) respectively (scores: 0.23 and 0.20). This provides additional support for the evolutionary relationship between these cell types, consistent with the recent report hahnEvolutionNeuronalCell2023a . The correspondence is particularly compelling as it respects both response polarity (ON vs. OFF) and kinetics (sustained vs. transient), suggesting conservation of fundamental visual processing channels across mammalian evolution.
Importantly, OT-MESH identified the strongest overall correspondence between macaque fRGC12 and mouse type C10 (an ON direction-selective RGC) with a high alignment score of 0.66 (SupplementaryTable 2). This finding suggests a previously underappreciated evolutionary relationship between these types. Notably, the existence of an ON direction-selective RGC in primate retina was independently confirmed by Wang et al. wangONtypeDirectionselectiveGanglion2023 , lending strong external validation to our method’s predictive power. We also detected robust correspondences between macaque fRGC14 and mouse melanopsin-expressing ipRGCs (C40 and C33, corresponding to M1 subtypes) (scores: 0.60 and 0.40), confirming the conservation of this non-image-forming visual pathway across mammals; and the mapping between macaque fRGC11 and mouse F-mini-ON cells (C3) (score: 0.22), suggesting potential evolutionary links between these functionally specialized types. The correspondence matrix also revealed that many mouse RGC types lacked clear macaque counterparts, suggesting evolutionary specialization in the murine lineage.
@@ 6 Discussion
We have presented OT-MESH, an unsupervised computational framework that systematically identifies evolutionary correspondences between cell types across species. The core innovation lies in the integration of the MESH procedure with an SNR-based gene selection and centroid-based OT cost matrix. This combination transforms standard OT outputs into sparse and biologically interpretable correspondence matrices. Applied to retinal neuron datasets from mouse and macaque, OT-MESH successfully recovered known homologies and revealed novel, biologically plausible correspondences. Its ability to quantify alignment strength and robustly handle differences in cell type numbers enhances its utility for cross-species comparative analysis. The observed patterns of conservation (e.g., in BCs) and divergence (e.g., more specialized RGCs in mouse) align with existing biological understanding, while newly predicted matchings offer clear avenues for future experimental validation.
Compared to existing methods, OT-MESH demonstrates several clear advantages. In contrast to reference-based classifiers such as XGBoost pengMolecularClassificationComparative2019a ; zhangEvolutionaryDevelopmentalSpecialization2024 ; wangMolecularCharacterizationSea2024 , OT-MESH is unsupervised and symmetric, requiring no designation of reference versus query species. By operating directly on cell-type centroids rather than individual cells, it achieves significant computational efficiency compared to cell-level projection-based matching methods such as Harmony korsunskyFastSensitiveAccurate2019 . Unlike standard entropy-regularized OT, which yields overly diffuse and uninterpretable transport plans, the MESH refinement produces sparse correspondence matrices that align closely with biological intuition and ground truth. This interpretability enables direct hypothesis generation, as exemplified by the high-confidence alignment between macaque fRGC12 and mouse ON direction-selective RGC (C10), which was independently validated by Wang et al. wangONtypeDirectionselectiveGanglion2023 . This finding illustrates how OT-MESH not only recovers known relationships but also predicts new ones with potential biological relevance.
While this study focused on transcriptomic data, our framework is readily extensible. Future versions could incorporate additional modalities, such as epigenomic states, spatial localization, or electrophysiological properties, to more robustly define cell types and resolve ambiguities that transcriptomics alone may not disentangle. Similarly, although SNR-based gene selection and cosine distance proved effective and computationally tractable, alternative feature selection strategies or dissimilarity metrics beyond SNR and cosine distance could also be explored. Beyond the retina, OT-MESH can be applied to explore evolutionary mapping of cell types across other organ systems (e.g., cortical neurons, immune cell lineages) and more distant species.
In conclusion, by enabling unsupervised, interpretable, and biologically grounded cell-type correspondence, OT-MESH can advance the field’s ability to trace the evolution of cellular diversity and specialization across species, offering a valuable tool for comparative genomics and evolutionary biology.
@@
fRGC12 C10 ON direction-selective 0.6616
PG_OFF C45 OFF transient alpha 0.6060
fRGC14 C40 M1 ipRGC 0.5967
MG_ON C43 ON sustained alpha 0.5242
fRGC5 C8 0.4058
fRGC14 C33 M1 ipRGC 0.3971
PG_OFF C34 0.3524
PG_ON C41 ON transient alpha 0.2345
fRGC11 C3 F-mini-ON 0.2156
MG_OFF C42 OFF sustained alpha 0.1995

### v2

@@ 5.2.2 Cross-Species Correspondence of Bipolar Cell (BC) Types
We next applied OT-MESH to cross-species alignment of BC types between mouse (15 types) and macaque (12 types). The resulting correspondence matrix (Figure 4A) and matched cell type graph (SupplementaryFigure 3A) reveal clear evolutionary matches, forming strong diagonal-like patterns.
@@
Alignment scores (SupplementaryTable 2) quantify these correspondences. The strongest match was between mouse RBC and macaque RB (rod BCs), with a score close to 1. Among cone BCs, macaque DB1, DB2, DB3a, DB3b, DB6, IMB, and BB/GB* types were uniquely and strongly assigned to mouse BC2, BC4, BC3A, BC3B, BC6, BC7, and BC8/9 types, respectively, consistent with previous reports hahnEvolutionNeuronalCell2023a ; shekharCOMPREHENSIVECLASSIFICATIONRETINAL2016a .
@@
5.2.3 Cross-Species Correspondence of Ganglion Cell (RGC) Types
@@
Most notably, our analysis confirmed the correspondence between mouse alpha RGCs and primate midget/parasol ganglion cells with high confidence (SupplementaryTable 3, SupplementaryFigure 3B). Specifically, mouse OFF transient alpha (C45) maps strongly to macaque OFF parasol ganglion cells (PG-OFF) (score: 0.71), and ON sustained alpha (C43) corresponds to macaque ON midget ganglion cells (MG-ON) (score: 0.72). Similarly, mouse ON transient alpha (C41) and OFF sustained alpha (C42) map to macaque ON parasol (PG-ON) and OFF midget ganglion cells (MG-OFF) respectively (scores: 0.67 and 0.30). This provides additional support for the evolutionary relationship between these cell types, consistent with the recent report hahnEvolutionNeuronalCell2023a . The correspondence is particularly compelling as it respects both response polarity (ON vs. OFF) and kinetics (sustained vs. transient), suggesting conservation of fundamental visual processing channels across mammalian evolution.
Importantly, OT-MESH identified a strong correspondence between macaque fRGC12 and mouse type C10 (an ON direction-selective RGC) with a high alignment score of 0.71 (SupplementaryTable 3, SupplementaryFigure 3B). This finding suggests a previously underappreciated evolutionary relationship between these types. Notably, the existence of an ON direction-selective RGC in primate retina was independently confirmed by Wang et al. wangONtypeDirectionselectiveGanglion2023 , lending strong external validation to our method’s predictive power. We also detected robust correspondences between macaque fRGC14 and mouse melanopsin-expressing ipRGCs (C40 and C33, corresponding to M1 subtypes) (scores: 0.70 and 0.68), confirming the conservation of this non-image-forming visual pathway across mammals; the mapping between macaque fRGC7 and a mouse ON orientation-selective RGC (C27) (score: 0.67), and the mapping between macaque fRGC11 and mouse F-mini-ON cells (C3) (score: 0.67), suggesting potential evolutionary links between these functionally specialized types. The correspondence matrix also revealed that many mouse RGC types lacked clear macaque counterparts, suggesting evolutionary specialization in the murine lineage.
@@ 6.1 Summary of Key Findings
We have presented OT-MESH, an unsupervised computational framework that systematically identifies evolutionary correspondences between cell types across species. The core innovation lies in the integration of the MESH procedure with SNR-based gene selection and centroid-based OT cost matrix construction. This combination transforms standard OT outputs—which are typically diffuse and difficult to interpret—into sparse and biologically meaningful correspondence matrices. Applied to retinal neuron datasets from mouse and macaque, OT-MESH successfully recovered known homologies and revealed novel correspondences, including the prediction of an ON direction-selective RGC in primates that was subsequently validated experimentally wangONtypeDirectionselectiveGanglion2023 . The observed patterns of conservation in BCs and greater divergence in RGCs align with current understanding of retinal evolution, while maintaining computational efficiency crucial for practical applications.
6.2 Methodological Advantages
OT-MESH demonstrates several key advantages over existing approaches. Unlike reference-based methods such as XGBoost pengMolecularClassificationComparative2019a ; zhangEvolutionaryDevelopmentalSpecialization2024 , OT-MESH is unsupervised and symmetric, eliminating bias from reference species designation. By operating on cell-type centroids rather than individual cells, it achieves orders of magnitude better computational efficiency than projection-based methods like Harmony korsunskyFastSensitiveAccurate2019 . Most critically, the MESH refinement procedure addresses the fundamental limitation of standard entropy-regularized OT by actively promoting sparsity through iterative entropy minimization. This results in correspondence matrices that are both mathematically optimal and biologically interpretable.
Our synthetic experiments demonstrated that OT-MESH achieves performance comparable to RefCM-Strict—which has perfect knowledge of the correspondence structure—while being 54× faster and not requiring any prior constraints. This dramatic speedup stems from fundamental algorithmic differences: RefCM computes exact EMD, which has complexity $O(n^{3}log(n))$ for n cell types. In contrast, OT-MESH leverages entropy-regularized optimal transport solved via the Sinkhorn algorithm, whose complexity has been proven to be nearly $O(n^{2})$. For the large-scale atlases emerging from current consortium efforts—often containing hundreds to thousands of cell types—this complexity difference becomes prohibitive. Combined with OT-MESH’s robustness to noise (maintaining ARI > 0.93 even at 8× noise levels), these computational advantages make OT-MESH particularly suitable for systematic comparative analyses across multiple species, organs, or developmental stages.
6.3 Limitations and Future Directions
6.3.1 Dependence on Upstream Analysis
While our noise robustness experiments demonstrate that OT-MESH maintains high performance even under challenging conditions (ARI > 0.93 at 8× noise), the method still depends on the quality of upstream analyses. High-quality gene orthology mapping is essential for accurate cross-species comparisons, and errors in ortholog assignment could propagate through the analysis. Similarly, the method assumes that input cell types have been appropriately identified through clustering or other means. However, our results suggest that OT-MESH is remarkably robust to technical noise. Future work could explore integration with simultaneous clustering and alignment methods to reduce dependence on pre-defined cell type annotations.
6.3.2 Uniform Marginals and Unbalanced Transport
Our current implementation assumes uniform marginal distributions, reflecting an agnostic prior about cell type correspondences. While this works well for the retinal datasets studied here, it may be suboptimal for scenarios with many species-specific cell types or when prior knowledge suggests unequal correspondence probabilities. In cases where one species has evolved numerous specialized cell types absent in the other, the uniform marginal assumption could force spurious matches. Extensions using unbalanced optimal transport chizatScalingAlgorithmsUnbalanced2017 , which allows for creation and destruction of mass, could better handle such scenarios by explicitly modeling unmatched types. This would be particularly valuable for comparisons across evolutionarily distant species or for organs undergoing rapid evolutionary change.
6.3.3 Choice of Features and Distance Metrics
We chose SNR for gene selection because it specifically identifies genes that distinguish cell types (high between-type variance) from those with high technical noise (high within-type variance), making it well-suited for building robust cell-type centroids. Similarly, cosine distance was selected for its insensitivity to expression magnitude differences that commonly arise from technical or biological variation between species. However, these choices may not be optimal for all datasets. For instance, datasets with strong batch effects might benefit from more sophisticated feature selection methods, while comparisons focusing on absolute expression levels might require alternative distance metrics. Future work could explore adaptive selection of these components based on dataset characteristics, or ensemble approaches that combine multiple feature sets and metrics.
6.3.4 Extension to Multiple Modalities
We only tested our methodology on transcriptomic data, but cell type identity is determined by multiple molecular and functional properties. Future versions could incorporate chromatin accessibility, DNA methylation, spatial location, morphology, or electrophysiological features. Multi-modal integration could resolve ambiguities in purely transcriptomic comparisons and provide more comprehensive evolutionary insights. The OT framework naturally extends to such settings through appropriate cost matrix design that combines multiple data types.
7 Conclusion
OT-MESH provides a principled, efficient, and interpretable solution to the fundamental problem of cross-species cell type matching. By combining optimal transport theory with entropy minimization, it achieves the accuracy of constrained methods while maintaining the flexibility and efficiency needed for exploratory analysis. As single-cell atlases expand across the tree of life, tools like OT-MESH will be essential for synthesizing these data into a unified understanding of cellular evolution. The framework’s extensibility to unbalanced transport, multiple modalities, and hierarchical matching positions it well for future developments in comparative genomics and evolutionary cell biology.
@@ 7 Conclusion
MG_ON C43 ON sustained alpha 0.7236
PG_OFF C45 OFF transient alpha 0.7137
fRGC12 C10 ON direction-selective 0.7079
PG_OFF C34 0.7064
fRGC14 C40 M1 ipRGC 0.7001
fRGC5 C8 0.6938
fRGC14 C33 M1 ipRGC 0.6801
fRGC7 C27 ON orientation-selective 0.6740
PG_ON C41 ON transient alpha 0.6735
fRGC11 C3 F-mini-ON 0.6714